# Score-based Generative Priors Guided Model-driven Network for MRI Reconstruction

### v1

@@ IV-A Score matching and Langevin dynamics
where $\xi_{t}\sim\mathcal{N}(0,\textit{I})$ and $t=0,1,2,..,T$. $\eta_{t}$ is the step size in the $t$th iteration and $\eta_{t}=\epsilon\cdot\frac{\sigma^{2}_{j}}{\sigma^{2}_{L}}$. $\{\sigma_{j}\}_{j=1}^{L}$ is a sequence of noise scales that gradually reduces each step size in the annealed Langevin dynamics, where $\sigma_{1}>\sigma_{2}>\cdots>\sigma_{L}$. $\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{t})}{\gamma_{t}^{2}+\sigma^{2}}$ is the consistency term introduced to exploit known measurements. Finally, $x_{T}$ approaches the target as $T$ becomes sufficiently large, which is used as a guide in the follow-up steps.
IV-B Denoising module
Over-smoothing details and hallucination artifacts were observed on the PGIs obtained in the first step, and the denoising step aimed to initially eliminate artifacts and compensate for missing details. First, we designed a KI-branch to learn an denoising mapping from the data between the frequency domain and the image domain. Considering that artifacts usually have different frequencies, which makes it possible to separate them in the k-space, we transferred $x_{T}$ back to the k-space and used a U-Net with a skip connection to learn the mapping to the denoised outputs. Subsequently, U-Net learned to eliminate the noisy signals in k-space. A data consistency (DC) block was concatenated to the U-Net to maintain consistency with the measurements. In addition, the complementarity between the k-space and image domain could also help obtain better performance. We then used an inverse Fourier transform to transfer the k-space data back to the image domain, which is denoted as $x_{d}$, as shown in Fig. 2 (b).
Then, we designed an input-enriched U-Net, marked as $N_{d}$, to further eliminate the noises in $x_{d}$. Fundamentally, the score network is used to generate clean images from noises and the artifacts are generated during the sampling process. Consequently, there could exist an underlying mapping between the artifacts and the elements involved in the sampling process (i.e. equation (8)). Therefore, with the input $x_{T}$ in the DM, we fed $x_{T}$, $s_{\theta}(x_{T};\sigma_{T}$) and $\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{T})}{\gamma_{T}^{2}+\sigma^{2}}$ as the input elements. We fine-tuned $s_{\theta}$, marked as $f_{s_{\theta}}$ that we removed the last batch normalization and activation layers to avoid undesirable loss of information, and concatenated a fine-tuning convolutional (FTC) layer to learn the noise pattern from the features. Only the FTC layer in $f_{s_{\theta}}$ was updated during training. The parameters in the other parts of $f_{s_{\theta}}$ were fitted, and the Langevin sampling results were fed to $f_{s_{\theta}}$ such that the FTC layers could directly learn the noise pattern from the score-matching process for each image.
@@ IV-B Denoising module
Besides, to avoid newly introduced noises when obtaining $x_{d}$ from $x_{T}$, $x_{d}$ was also fed to $N_{d}$. Therefore, the enriched input features of $N_{d}$ have four components. By concatenating and feeding the four components into the DM, the denoising process is denoted as:
$x_{T}^{0}=x_{d}+N_{d}\{x_{p},x_{T},\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{T})}{\gamma_{T}^{2}+\sigma_{T}^{2}},x_{d}\}$ (9)
@@
Figure 3: Examples of ground truth images (the first column), undersampled images (the second column) and the corresponding k-space trajectories (the third column). The images from the first and second rows are selected from the Knee dataset with four-fold and six-fold random Cartesian mask. The images from the third and fourth row are selected from the fastMRI dataset with four-fold and six-fold equispaced Cartesian mask.
coronal PD 4$\times$ acceleration
coronal PD 6$\times$ acceleration
FastMRI 4$\times$ acceleration
FastMRI 6$\times$ acceleration
Figure 4: Examples of the reconstructed images of different methods from coronal PD sequence and FastMRI dataset at 4$\times$ and 6$\times$ acceleration. The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps corresponding to the objective reality. The methods from left to right: 1. objective reality; 2. zero-filled; 3. TV; 4. U-Net; 5. D5C5; 6. ISTA-Net; 7. VS-Net; 8. E2EVarNet; 9. ReVarNet; 10. MeDL-Net; 11. SGM-Net (ours).
IV-C Unrolled network with PGI-guided densely-connected cascades
The third step of the proposed workflow is to design an unrolled network to obtain higher-quality reconstructions from denoised guidance images (DGIs) (that is, $x_{T}^{0}$), and the undersampled measurements $x_{z}^{0}$ and $x_{T}^{0}$ are used as two different constraints during unrolled network training. As shown in Fig. 2, the network consists of several DGI-guided densely connected cascades (GICs). The inputs of GICs are $x_{T}^{k}$ and $x_{z}^{k}$ where $k=0,1,...,K$, and $x_{z}^{0}$ is the coil-combined zero-filled image of the undersampled measurements. In each GIC, the undersampled measurements $x_{z}^{k}$ were periodically updated to approach the target images, whereas the DGIs $x_{T}^{k}$ were simultaneously updated to provide more accurate guidance, forming a parallel network structure. $I$ Data fidelity (DF) and $I$ U-Net blocks were alternately stacked in each GIC to solve (7), and U-Net blocks were used to learn the gradient of $\Psi(x)$ from the data. For the $x_{T}^{k}$ branch, the calculations in DF blocks can be represented as
@@ IV-C Unrolled network with PGI-guided densely-connected cascades
where $x_{z}^{k_{0}}$ is equal to $x_{z}^{k}$. In addition, we propose a densely connected structure to enrich the input information for each U-Net block. For the $x_{T}^{k}$-branch, the initial input and each output of DF blocks corresponding to a U-Net block are concatenated before being fed into U-Net. Therefore, we have:
$x_{T}^{k_{i+1}}=U_{T}^{i}(\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}}\})$ (12)
where $U_{T}^{i}$ denotes the $i$th U-Net block in the $x_{T}^{k}$-branch, $x_{T}^{k}$ denotes the input to the corresponding branch, and $r_{T}^{k_{i}}$ denotes the output of the $i$th DF block. A similar strategy is implemented in the $x_{z}^{k}$-branch. Considering that we take the final output of the $x_{z}^{k}$ branch in the last GIC as the final reconstruction, the $x_{z}^{k}$ branch is essentially the main branch, and the $x_{T}^{k}$ branch is simultaneously used to guide the reconstruction in the main branch. Therefore, the concatenated information from the $x_{T}^{k}$ branch is also fed to the $x_{z}^{k}$ branch to learn from the features from different stages, represented as
$x_{z}^{k_{i+1}}=U_{z}^{i}(\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{i+1}}\})$ (13)
@@
$x_{T}^{k+1}=DC(x_{T}^{k}+x_{T}^{k_{I}})$ (14)
For $x_{z}^{k}$, we further adopted a four-layer attention module (AM) [64] between the last U-Net and skip connection. The information fed to the last DF block is concatenated and simultaneously fed to the AM to produce an attention map
$M_{A}=AM(\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{i+1}}\})$ (15)
@@
$x_{z}^{k+1}=DC(x_{z}^{k}+x_{z}^{k_{I}}\odot M_{A}+x_{T}^{k_{I}}\odot((\textbf{J}-M_{A}))$ (16)
@@
where J is an all-ones matrix and $\odot$ denotes the Hadamard product. Element-wise attention results in a deeper fusion between the two branches and optimizes the intermediate results in each $x_{z}^{k}$ branch.
IV-D loss functions
In this study, the loss function for unrolled network training was designed as
$Loss=\gamma_{1}(x_{d}+x_{T}^{0}+x_{z}^{K})+\gamma_{2}(\sum_{k=1}^{K}x_{T}^{k}+\sum_{k=1}^{K}x_{z}^{k})$ (17)
where $x_{d}$ amd $x_{T}^{0}$ are important guidelines for ensuring reconstruction performance, and $x_{z}^{K}$ is the result. $\sum_{k=1}^{K}x_{T}^{k}$ and $\sum_{k=1}^{K}x_{z}^{k}$ are intermediate results. We set $\gamma_{1}=1$ and $\gamma_{2}=0.1$ to distribute more weights for the guidance and results and fewer weights to help the network converge.
IV-E Training details
@@
We adopted the score network $s_{\theta}$ released by the authors in [25], which is pretrained on a brain dataset. We followed the proposed hyperparameters in that manuscript [^1] https://github.com/utcsilab/csgm-mri-langevin during sampling, and we do not retrain the network on our knee dataset. The numbers of the four-cascade filters were 128, 256, 512, and 1024. The number of filters in the FTC layer is set to 2. In SGM-Net, $K$ is set to five and $I$ is set to three. The numbers of filters in all U-Nets in the DM and GICs were set to 16, 32, 64, and 128. The step size in DF blocks was learned directly in the network. All complex-valued MR images were stored in two channels during the network training. In particular, the first U-Net in DM uses k-space data as input, and the coil channel is combined with the feature channel. Therefore, the number of input and output channels of U-Net was set to 30 (15 coils). The proposed SGM-Net was trained on a UBUNTU 20.04 LTS system with a 3090 GPU (24GB) for 100 epoches, and the learning rate was set to 1e-3. The batch size is set to 1.
V Experiments and Results
V-A Dataset
@@
TABLE I: The quantitative results on Coronal PD and PDFS sequences, as well as the FastMRI dataset, compared with those of cutting-edge methods.
@@
SGM-Net (ours) 38.9212$\pm$3.4296 37.9430$\pm$3.2521 0.9344$\pm$0.0604 0.9234$\pm$0.0067
In this study, we adopted a publicly available 15-coil knee dataset proposed in [51] to train our model and compared it with other methods. The knee dataset consisted of five 2D turbo spin-echo (TSE) sequences, each of which contained volumes scanned using a 3T scanner from 20 patients. We conducted experiments on coronal proton density (PD; TR=2750 ms, TE=27 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 35-42 slices), and coronal fat-saturated PD (PDFS; TR=2870 ms, TE=33 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 33-44 slices) sequences. The coil sensitivity maps were precomputed using ESPRIRiT and included in the dataset. The training, validation, and testing sets were randomly divided at a ratio of 3:1:1. Random Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration. 8% lines at the center region are preserved. Examples of training images and the corresponding k-space trajectories are shown in Fig. 3 (the first and second rows).
FastMRI [8] is a publicly available dataset for accelerated MRI, which has been adopted in relevant studies. Considering the limitations of our hardware, we randomly selected a subset of the 15-coil knee dataset to evaluate our model and demonstrate its effectiveness. The data were acquired using the 2D TSE protocol, containing PD and PDFS sequence scanning from 3T or 1.5T systems (TR=2200-3000 ms, TE=27-34 ms, in-plane resolution 0.5$\times$0.5 mm2, slice thickness 3 mm). A total of 419, 33, and 375 slices were used for training, validation, and testing sets, respectively. Coil sensitivity maps were not available in the dataset, and related studies have been conducted to estimate and optimize sensitivity maps during training. However, we did not focus on SM estimation in this study and used pretrained SMs in our model to highlight the improvements brought about by the proposed methods. In the fastMRI dataset, we calculated the SMs using the sigpy [65] toolbox. Equispaced Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration. 8% lines at the center region are preserved. Examples of training images and the corresponding k-space trajectories are shown in Fig. 3 (the third and fourth rows).
V-B comparisons with the cutting-edge methods
@@
We compared the proposed SGM-Net with two traditional and seven DL-based methods, as presented in Table I. In the zero-fill method, the missing data in the measurements were filled with zero. The TV method uses total variation as the image prior in a CS-based approach. For DL-based methods (i.e., U-Net [8] [^2] https://github.com/facebookresearch/fastMRI, D5C5 [66] [^3] https://github.com/js3611/Deep-MRI-Reconstruction. ISTA-Net [50] [^4] https://github.com/jianzhangcs/ISTA-Net-PyTorch VSNet [52] [^5] https://github.com/j-duan/VS-Net. E2EVarNet [53] [^6] https://github.com/NKI-AI/direct, ReVarNet [67]6 and MeDL-Net [68] [^7] https://github.com/joexy312/MEDL-Net), we adopted the released codes and retrained the networks on our device. We made a concerted effort to keep the announced settings same as those in the corresponding articles, except that we reduced the number of filters in ReVarNet so that all networks could be trained on one GPU. All methods were implemented on the knee dataset (coronal PD and PDFS sequences) and a subset of fastMRI dataset. The PSNR and SSIM results were calculated under four- and six-fold acceleration to evaluate the different methods. As shown in Table I, the proposed SGM-Net robustly achieves the best results on all the datasets with different acceleration factors. The reconstruction results of the sample images are presented in Fig.4. We can observe from the zoomed region (second rows) that our method recovered the most details, and the error maps (third row) exhibited the least errors.
V-C Ablation Studies

### v2

@@ 4.1 Score matching and Langevin dynamics
where $\xi_{t}\sim\mathcal{N}(0,\textit{I})$ and $t=0,1,2,..,T-1$. $\eta_{t}$ is the step size in the $t$th iteration and $\eta_{t}=\epsilon\cdot\frac{\sigma^{2}_{j}}{\sigma^{2}_{L}}$. $\{\sigma_{j}\}_{j=1}^{L}$ is a sequence of noise scales that gradually reduces each step size in the annealed Langevin dynamics, where $\sigma_{1}>\sigma_{2}>\cdots>\sigma_{L}$. $\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{t})}{\gamma_{t}^{2}+\sigma^{2}}$ is the consistency term introduced to exploit known measurements. Finally, $x_{T}$ approaches the target as $T$ becomes sufficiently large, which is used as a guidance in the follow-up steps.
4.2 Denoising module
The PGIs generated in the first step exhibits over-smooth details and hallucination artifacts, and the denoising step aims to coarsely eliminate the noises and artifacts in PGIs. To achieve this, we designed a CIE and a SIE block to extract features, which were then used within a residual structure to generate denoised outputs. Consequently, the mappings between extracted features and the noise patterns were established.
The CIE blocks were designed to exploit the complementarity between the frequency domain and the image domain. As depicted in Fig 2 (b), first we used the expand operator to obtain the corresponding k-space data of $x_{T}$, marked as $y_{T}$. Then, $y_{T}$ was reshaped and the coil channel was combined with the feature channel before fed to a U-Net. In this study, all complex-valued MR images were stored in two distinct channels during the network training. Therefore, the number of input and output channels of the U-Net was set to 30 (with 15-coil data). A data consistency (DC) block was concatenated to the output of U-Net, marked as $y_{c}$, to maintain consistency with the measurements, and the $j$th coefficient in $y_{c}$ was updated as:
$y_{d}(j)=D(y_{c}(j))=\left\{\begin{array}[]{l}y_{c}(j),\ \ \ \ \ \ \ \ \ j\notin\Omega\\ \frac{y_{c}(j)+\mu y(j)}{1+\mu},\ \ \ \ j\in\Omega\\ \end{array}\right.$ (9)
where $\Omega$ is the subset of sampled points from the complete k-space data. $y_{c}$ and $y_{d}$ are the input and output of the DC blocks, respectively, and $y$ is the undersampled measurements. We then used an reduce operator to transfer the k-space data back to the image domain, which is denoted as $x_{d}$. The CIE blocks could extract the features associated with the noise patterns directly in k-space. Considering the other components in DMs operate in the image domain, the output $x_{d}$ is expected to compensate for the missing information in the other blocks, thereby enhancing overall performance.
Essentially, score network is used to generate clean images from noise, and the artifacts can be considered as generated during the sampling process. Consequently, there could exist an underlying mapping between the artifacts and the elements involved in the score network. Therefore, we designed a SIE to extract features correlated to the noise patterns from the PSN. As depicted in Fig 2 (b), the PSN is a six-layer RefineNet and the features in deeper RefineNet blocks (RBs) closely associated with the score of the inputs, which could potentially compromise noise-related information. Moreover, considering that the optimal level of features for learning the noise is unrevealed, we fine-tuned the PSN and extract features across multiple layers. Specifically, to build the mapping to the noise in $x_{T}$, we directly fed $x_{T}$ to PSN and during training, all the parameters in the PSN was fitted. Then, we concatenated convolutional blocks to the last four RBs in PSN to learn multi-level features. Each convolutions was followed by a batch normalization and a ReLU function. Then, deconvolutions were used to double the feature sizes, and the features were concatenated with the counterparts extracted in the shallower RBs. Each convolutions and deconvolutions was followed by a batch normalization block and a ReLU function. $1\times 1$ convolutions were first applied to the output of RBs to adjust the number of channels, thereby constraining the module scale. As a results, different levels of features can be simultaneously captured to maximize the stability and accuracy of the mapping established to noise.
@@ 4.2 Denoising module
The output of the SIE is marked as $x_{p}$, which was concatenated with $x_{d}$ and fed to a U-Net, marked as $N_{d}$, to learn the noise in $x_{T}$. Besides, to avoid newly introduced noises when obtaining $x_{d}$ and $x_{p}$ from $x_{T}$, $x_{T}$ was also fed to $N_{d}$. Therefore, the enriched input features of $N_{d}$ have three components. By concatenating and feeding them into the $N_{d}$, the denoising process is denoted as:
$x_{T}^{0}=x_{T}+N_{d}\{x_{T},x_{p},x_{d}\}$ (10)
@@
4.3 Unrolled network with PGI-guided densely-connected cascades
The DM aims to realize a coarse denoising, and the third step of the proposed workflow is to further improve the image quality and recover the fine details from denoised guidance images (DGIs) (that is, $x_{T}^{0}$). The undersampled measurements $x_{z}^{0}$ and $x_{T}^{0}$ were used as two different known information during unrolled network training. As shown in Fig. 2, the network consists of several DGI-guided densely connected cascades (GICs). The inputs of the $(k+1)$th GICs are $x_{T}^{k}$ and $x_{z}^{k}$ where $k=0,1,...,K-1$, and $x_{z}^{0}$ is the coil-combined zero-filled image of the undersampled measurements. In each GIC, $x_{z}^{k}$ were periodically updated to approach the target images, whereas the DGIs $x_{T}^{k}$ were simultaneously updated to provide more accurate guidance, forming a parallel network structure. $I$ Data fidelity (DF) and $I$ U-Net blocks were alternately stacked in each GIC to solve (7), and U-Net blocks served as learning based regularizers. For the $x_{T}^{k}$ branch, the calculations in DF blocks can be represented as
@@ 4.3 Unrolled network with PGI-guided densely-connected cascades
where $x_{z}^{k_{0}}$ is equal to $x_{z}^{k}$. In addition, we proposed a densely connected structure to enrich the input information for each U-Net block. This is based on the idea that DL-based methods are more flexible than the conventional algorithms, allowing for learning from richer inputs across various stages and extracting more valuable features. Consequently, the densely connected information can mitigate the impact of inaccurate cascades and establish robust mappings to the targets. For the $x_{T}^{k}$-branch, the initial input and each output of DF blocks corresponding to a U-Net block were concatenated before being fed into U-Net. Therefore, we have:
$x_{T}^{k_{i+1}}=U_{T}^{i+1}\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}}\},i=0,1,...,I-1$ (13)
where $U_{T}^{i+1}$ denotes the $(i+1)$th U-Net block in the $x_{T}^{k}$-branch, $x_{T}^{k}$ denotes the input to the corresponding branch, and $r_{T}^{k_{i}}$ denotes the output of the $i$th DF block. A similar strategy was implemented in the $x_{z}^{k}$-branch. Considering that we take the final output of the $x_{z}^{k}$ branch in the last GIC as the final reconstruction, the $x_{z}^{k}$ branch is essentially the main branch, and the $x_{T}^{k}$ branch is simultaneously used to guide the reconstruction in the main branch. Therefore, the concatenated information from the $x_{T}^{k}$ branch was also fed to the $x_{z}^{k}$ branch to learn from the features in different stages, represented as
$x_{z}^{k_{i+1}}=U_{z}^{i+1}\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{i+1}}\},i=0,1,...,I-1$ (14)
@@
$x_{T}^{k+1}=D(x_{T}^{k}+x_{T}^{k_{I}})$ (15)
For $x_{z}^{k}$, we further incorporated a four-layer attention module (AM), marked as $\mathcal{M}_{A}$, between the last U-Net and skip connection, as shown in Fig 2 (c). Information fed to the last U-Nets in the two branches were concatenated and simultaneously fed to the AM to produce an attention map
$m_{A}=\mathcal{M}_{A}\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{I}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{I}}\}$ (16)
@@
$x_{z}^{k+1}=D(x_{z}^{k}+x_{z}^{k_{I}}\odot m_{A}+x_{T}^{k_{I}}\odot((\textbf{J}-m_{A}))$ (17)
@@
where J is an all-ones matrix and $\odot$ denotes the Hadamard product. The element-wise attention results in a deeper fusion between the two branches and optimizes the intermediate results in each $x_{z}^{k}$ branch, which amplifies the likelihood for subsequent GICs to converge towards the targets.
4.4 loss functions
In this study, the overall loss for network training was designed as:
$\mathcal{L}{(x_{T}^{0},x_{g})}+\mathcal{L}{(x_{z}^{K},x_{g})}+\sum_{k=1}^{K}w_{k}[\mathcal{L}{(x_{T}^{k},x_{g})}+\mathcal{L}{(x_{z}^{k},x_{g})}]$ (18)
where $x_{g}$ is the target image. $x_{T}^{0}$ amd $x_{z}^{K}$ are the final outputs of DM and the unrolled network. $x_{T}^{k}$ and $x_{z}^{k}$ are intermediate results from the two branches in the unrolled network. $\{w_{k}\}_{k=1}^{K}$ is a sequence assigns heavier weights to the intermediate results from deeper GICs and $w_{k}=10^{\frac{k-K}{K-1}}$ [67]. When training on the coronal PD sequence and FastMRI dataset, we adopted mean square error as the loss function that $\mathcal{L}:=\mathcal{L}_{MSE}$. For the coronal PDFS sequence, we further included a SSIM loss that $\mathcal{L}:=\mathcal{L}_{MSE}+\mathcal{L}_{SSIM}$ which equally contribute with the MSE loss during training to improve reconstruction outcomes.
4.5 Implementation details
@@
We adopted the score network $s_{\theta}$ released by the authors in [24], which was pretrained on FastMRI brain dataset. We followed the proposed hyperparameters in that manuscript [^1] https://github.com/utcsilab/csgm-mri-langevin during sampling, and we did not retrain the network on our knee dataset. In SGM-Net, $K$ was set to five and $I$ was set to three. The numbers of filters in the U-Nets from DMs were set to 32, 64, 128, and 256. The numbers of filters in all U-Nets from GICs were set to 16, 32, 64, and 128, respectively. In SIE blocks, the number of output channels of the RBs were set to 128, 128, 256, 256, 256 and 512. The number of channels of the inputs and outputs in $1\times 1$ convolutions, which are concatenated to RBs, were set to (128, 16), (128, 32), (256, 64) and (256, 128). For the last $1\times 1$ convolution, the number of input and output channels were 16 and 2, respectively. The CBRs and DBRs halved the number of channels of their inputs. The step size in DF blocks and the $\mu$s in DC blocks were learned directly in the network. The proposed SGM-Net was trained on a UBUNTU 20.04 LTS system with a 3090 GPU (24GB) for 100 epoches, and the learning rate was set to 1e-3. The batch size was set to 1.
5 Experiments and Results
5.1 Dataset
Figure 3: Examples of ground truth images (the first column), undersampled images (the second column) and the corresponding k-space trajectories (the third column) from different datasets.
In this study, we adopted a publicly available 15-coil NYU knee dataset [53] to train our model and compared it with other methods. The knee dataset consisted of five 2D turbo spin-echo (TSE) sequences, each of which contained volumes scanned using a 3T scanner from 20 patients. Each volume consists of approximately 40 slices. We conducted experiments on coronal proton density (PD; TR=2750 ms, TE=27 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 35-42 slices), and coronal fat-saturated PD (PDFS; TR=2870 ms, TE=33 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 33-44 slices) sequences. The coil sensitivity maps were precomputed using ESPRIRiT and included in the dataset. The training, validation, and testing sets were randomly divided at a ratio of 3:1:1. Random Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration. 8% lines at the center region are preserved.
FastMRI [8] is a publicly available dataset for accelerated MRI, which has widely been adopted in relevant studies. Considering the limitations of our hardware, we randomly selected a subset of the 15-coil knee dataset to evaluate our model and demonstrate its effectiveness. The data were acquired using the 2D TSE protocol, containing PD and PDFS sequence scanning from 3T or 1.5T systems (TR=2200-3000 ms, TE=27-34 ms, in-plane resolution 0.5$\times$0.5 mm2, slice thickness 3 mm). A total of 419, 33, and 375 slices were used for training, validation, and testing sets, respectively. Coil sensitivity maps are not available in the dataset, and related studies have been conducted to estimate and optimize sensitivity maps during training. However, we did not focus on SM estimation in this study and used pretrained SMs in our model to highlight the improvements brought about by the proposed methods. In the FastMRI dataset, we calculated the SMs using the sigpy [69] toolbox. Equispaced Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration. 8% lines at the center region are preserved. Examples of training images and the corresponding k-space sampling trajectories are shown in Fig. 3.
5.2 comparisons with the cutting-edge methods
@@ 5.1 Dataset
Table 1: The quantitative results on Coronal PD and PDFS sequences, as well as the FastMRI dataset, compared with those of cutting-edge methods.
@@
vSHARP 38.0836$\pm$3.1283 35.7326$\pm$2.8239 0.9134$\pm$0.0645 0.8621$\pm$0.0917
SGM-Net (ours) 38.8241$\pm$3.3808 36.8730$\pm$3.0364 0.9344$\pm$0.0588 0.9094$\pm$0.0707
coronal PD 4$\times$ acceleration
coronal PD 6$\times$ acceleration
FastMRI 4$\times$ acceleration
FastMRI 6$\times$ acceleration
Figure 4: Examples of the reconstructed images of different methods from coronal PD sequence and FastMRI dataset at 4$\times$ and 6$\times$ acceleration. The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps corresponding to the ground truth images. The methods from left to right: 1. ground truth; 2. zero-filled; 3. TV; 4. U-Net; 5. D5C5; 6. ISTA-Net; 7. VS-Net; 8. E2EVarNet; 9. ReVarNet; 10. MeDL-Net; 11. vSHARP; 12. SGM-Net (ours).
@@
We compared the proposed SGM-Net with two traditional and eight DL-based methods, as presented in Table 1. In the zero-fill method, the missing data in the measurements were filled with zero. The TV method used total variation as the image prior in a CS-based approach. For DL-based methods (i.e., U-Net [8] [^2] https://github.com/facebookresearch/fastMRI, D5C5 [70] [^3] https://github.com/js3611/Deep-MRI-Reconstruction. ISTA-Net [52] [^4] https://github.com/jianzhangcs/ISTA-Net-PyTorch VSNet [54] [^5] https://github.com/j-duan/VS-Net. E2EVarNet [55] [^6] https://github.com/NKI-AI/direct, ReVarNet [71]6, MeDL-Net [72] [^7] https://github.com/joexy312/MEDL-Net and vSHARP [67]6), we adopted the released codes and retrained the networks on our device. We made a concerted effort to keep the announced settings same as those in the corresponding articles, except that we reduced the number of filters in ReVarNet so that all networks could be trained on one GPU. All methods were implemented on the knee dataset (coronal PD and PDFS sequences) and a subset of fastMRI dataset. The PSNR and SSIM results were calculated under four- and six-fold acceleration to evaluate different methods. As shown in Table 1, the proposed SGM-Net robustly achieves the best results on all the selected datasets with different acceleration factors. Examples of reconstruction images are presented in Fig.4. We can observe from the zoomed region (second rows) that our method recovered the most details, and the error maps (third row) exhibited the least errors.
5.3 Ablation Studies